Examine the convergence of the series


a) 3×4/5^2+5×6/7^2+7×8/9^2+.........


b) 1+4x+4^2x^2+4^3x^3+......(x>0)

1
Expert's answer
2022-01-06T12:19:50-0500

a)

"\dfrac{3\times4}{5^2}+\dfrac{5\times6}{7^2}+\dfrac{7\times8}{9^2}+..."

"=\displaystyle\sum_{n=1}^{\infin}\dfrac{(2n+1)(2n+2)}{(2n+3)^2}"

Use Test for Divergence


"a_n=\dfrac{(2n+1)(2n+2)}{(2n+3)^2}"

"\lim\limits_{n\to \infin}a_n=\lim\limits_{n\to \infin}\dfrac{(2n+1)(2n+2)}{(2n+3)^2}"

"=\lim\limits_{n\to \infin}\dfrac{(\dfrac{2n}{n}+\dfrac{1}{n})(\dfrac{2n}{2}+\dfrac{2}{n})}{(\dfrac{2n}{n}+\dfrac{3}{n})^2}"

"=\lim\limits_{n\to \infin}\dfrac{(2+\dfrac{1}{n})(2+\dfrac{2}{n})}{(2+\dfrac{3}{n})^2}=\dfrac{(2+0)(2+0)}{(2+0)^2}=1\not=0"

The series "\dfrac{3\times4}{5^2}+\dfrac{5\times6}{7^2}+\dfrac{7\times8}{9^2}+..." diverges by the Test for Divergence.


b)


"1+4x+4^2x^2+4^3x^3+......=\displaystyle\sum_{n=0}^{\infin}(4x)^n, x>0"

This is a geometric series with "a=1" and "r=4x."

If "0<4x<1," the geometric series converges and its sum is


"S=\dfrac{1}{1-4x}."

If "4x\geq1," the geometric series diverges.

Then the series


"1+4x+4^2x^2+4^3x^3+......=\displaystyle\sum_{n=0}^{\infin}(4x)^n"

converges for "0<x<\dfrac{1}{4}" and diverges for "x\geq \dfrac{1}{4}."